Depreciation by Three Methods; Partial Years

Perdue Company purchased equipment on April 1 for $49,950. The equipment was expected to have a useful life of three years, or 5,400 operating hours, and a residual value of $1,350. The equipment was used for 1,000 hours during Year 1, 1,900 hours in Year 2, 1,600 hours in Year 3, and 900 hours in Year 4.

Required:

Determine the amount of depreciation expense for the years ended December 31, Year 1, Year 2, Year 3, and Year 4, by (a) the straight-line method, (b) units-of-output method, and (c) the double-declining-balance method.

Note: FOR DECLINING BALANCE ONLY, round the multiplier to four decimal places. Then round the answer for each year to the nearest whole dollar.

a. Straight-line method

Year Amount
Year 1 $
Year 2 $
Year 3 $
Year 4 $

b. Units-of-output method

Year Amount
Year 1 $
Year 2 $
Year 3 $
Year 4 $

c. Double-declining-balance method

Year Amount
Year 1 $
Year 2 $
Year 3 $
Year 4 $

I have finished almost all of this but year 4 on DDB method.
Answer #1

Depreciation as per double declining method

Year 1 = (49950 * 2/3 *9/12) = $24,975

Year 2 = (49950 - 24975) * 2/3 = $16650

Year 3 = (49950 - 24975 - 16650) * 2/3 = 5550

Year 4 = (49950 - 24975 - 16650 - 5550) * 2/3 = $1850